# Multimodal Large Language Models to Support Real-World Fact-Checking

### v1

@@ 3.3 Models
We use OpenAI’s API to collect responses from GPT-4V (gpt-4-1106-preview). We also experiment with open-source MLLMs, including LLaVA-1.5 (7b and 13b, Liu et al. 2023b, a), InstructBLIP (Flan-T5-xxl, Dai et al. 2023), and MiniGPT (v2, Chen et al. 2023). All parameters are set to the default values, with max_tokens=300 for GPT-4V. These models are notably representative and competitive in a multimodal setup.111We hereafter omit the version numbers of the open-source models, using LLaVA(7b), LLaVA(13b), MiniGPT, and InstructBLIP instead.
3.4 Datasets
Fauxtography |
Cosmos |
MOCHEG |
Post-4V |
True |
200 |
240 |
267 |
81 |
False |
200 |
240 |
237 |
105 |
Total |
400 |
480 |
504 |
186 |
Fauxtography
is a multimodal fact-checking dataset sourced from the websites Snopes and Reuters Zlatkova et al. (2019). Due to constraints posed by the GPT-4V API, such as decreased query speeds after exceeding a daily limit, we select a random subset of 400 entries, ensuring a balanced composition of 200 true and 200 false instances.
COSMOS
is a dataset built from Snopes and News Outlets, and tailored for training models to autonomously identify image and text pairs that are out of context Aneja et al. (2023), i.e. falsely paired. We randomly select 240 true and 240 false samples from the test split.
MOCHEG
Post-4V
To mitigate potential data leakage, we obtain new data from the Snopes based on articles published after the release of GPT-4V (Sept. 26, 2023). We adhere to the same filtering approach used for MOCHEG, resulting in 186 samples.
Details of all data samples used are presented in Table 1. Overall, the sample distribution is balanced, containing only two types of labels: True and False.222We exclude data labeled as Mixture, Unproven, and other labels that lack clear veracity. Additionally, we collected the URLs of the web pages corresponding to the claims and their fact-checking articles.
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
Image |
Model |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
| Yes |
GPT |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
LLa.(7b) |
337 |
6 |
1 |
56 |
53.4 |
46.3 |
449 |
7 |
4 |
20 |
52.0 |
50.0 |
409 |
6 |
0 |
89 |
54.2 |
44.8 |
157 |
1 |
3 |
25 |
42.0 |
37.1 |
LLa.(13b) |
286 |
98 |
3 |
13 |
54.4 |
52.3 |
331 |
134 |
13 |
2 |
69.7 |
67.7 |
363 |
112 |
28 |
1 |
61.5 |
58.1 |
116 |
65 |
2 |
3 |
56.3 |
55.9 |
MiniGPT |
55.5 |
62.1 |
54.3 |
55.9 |
Instruct. |
54.3 |
74.4 |
52.6 |
59.1 |
| No |
GPT |
105 |
202 |
32 |
61 |
80.1 |
71.0 |
111 |
253 |
50 |
66 |
64.0 |
57.5 |
153 |
265 |
38 |
48 |
73.9 |
68.9 |
35 |
68 |
10 |
73 |
74.8 |
67.7 |
LLa.(7b) |
398 |
0 |
2 |
0 |
50.3 |
50.0 |
474 |
0 |
4 |
2 |
50.2 |
49.8 |
502 |
0 |
2 |
0 |
53.0 |
52.8 |
181 |
0 |
4 |
1 |
43.1 |
43.0 |
LLa.(13b) |
167 |
214 |
17 |
2 |
62.5 |
62.0 |
263 |
195 |
20 |
2 |
76.6 |
75.8 |
168 |
316 |
20 |
0 |
62.0 |
61.5 |
40 |
124 |
22 |
0 |
60.4 |
60.8 |
MiniGPT |
56.8 |
60.4 |
54.0 |
58.1 |
Instruct. |
55.8 |
68.3 |
48.6 |
61.8 |
Yes |
Baseline |
50.0 |
50.0 |
53.0 |
56.5 |
4 Evaluation and Analysis
4.1 RQ1: Comprehensive Performance
Table 2 presents the comprehensive performance of various models across four datasets. The "Image" column clarifies the input type, with "Yes" indicating multimodal inputs and "No" signifying text-only inputs. MiniGPT and InstructBLIP, while incapable of providing explanations and confidence scores, still manage to generate predictions; thus, only the Overall Accuracy metric is assessed. The accuracy of the majority class is noted at the bottom of the table.
The upper section of Table 2 reveals that GPT-4V outperforms other models in both True/False and Overall Accuracy metrics, with True/False Accuracy frequently exceeding Overall Accuracy. It implies GPT-4V’s heightened precision when it responds with confidence. LLaVA’s accuracy consistently improves with increased model size. MiniGPT and InstructBLIP surpass the baseline and LLaVA(7b), nearing the accuracy levels of the 13b model. In addition, LLaVA(7b) shows a strong bias, almost always predicting True. As the model size increases, predictions become more balanced, and the expression of uncertainty rises. Furthermore, there is a significant reduction in the number of Deny responses, indicating that larger models are more capable of following instructions. The high Deny responses from GPT-4V are mainly due to its safety alignment, often replying with statements like "I’m sorry, but I cannot assist with this request." or "This content may violate our content policy."
Figure 2 displays multimodal claims and corresponding responses from GPT-4V and LLaVA(13b), with the annotations w. Image and w/o Image next to the model names indicating whether an image is in the input, respectively. From the perspective of explanations, GPT-4V’s responses contain a richer set of facts. For instance, it explicitly states, "there was never an official campaign for Michelle Obama in the 2020 election," in the 1st case. The 5th example provides detailed background information about Margaret Hamilton, including her role in "leading the development of onboard flight software for the Apollo mission." Moreover, GPT-4V adeptly incorporates other details from the images, such as referencing a "Google image search: Creepy Joe Biden" in the 4th image. In comparison, LLaVA(13b) lacks the capability to counteract manipulated images, as evidenced in the 2nd and 4th examples. Additionally, LLaVA’s explanations are typically brief, often limited to a description of the image followed by a conclusion about its veracity, indicating inefficient reasoning.
4.2 RQ2: Impact of Images: Multimodal VS. Text-Only inputs
By comparing data in Table 2, it is evident that images significantly impact the predictions of GPT-4V, with the model’s accuracy decreasing to various extents. Simultaneously, the instances where the model refuses to respond have noticeably increased. As in Figure 2, GPT-4V’s explanations indicate its awareness of missing images, using phrases like "there is no verified public record" or "without any provided context." This suggests that GPT-4V engages in knowledge retrieval to verify the existence or context of the images.
Counterintuitively, open-source models demonstrate higher accuracy with text-only inputs in many cases compared to multimodal inputs. To understand this, we scrutinize their explanations and find two specific scenarios. First, when the claim is false and the image is altered, as illustrated in the 1st and 4th examples in Figure 2, models can accurately predict the veracity of textual claims. However, when images are included, the models, failing to recognize the alterations, are misled into making incorrect predictions. Second, when both the image and the claim are true, as demonstrated in the 3rd example, models that consider images might deem the claim False due to inadequate evidence. In contrast, predictions for textual claims remain true, even with hallucinations highlighted in yellow. This implies that the models were not exposed to enough negative examples during the fine-tuning phase.
| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
EN |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
ZH |
144 |
188 |
48 |
20 |
78.6 |
75.3 |
180 |
168 |
100 |
32 |
83.0 |
72.5 |
216 |
198 |
50 |
40 |
84.3 |
79.2 |
62 |
82 |
30 |
12 |
78.5 |
68.3 |
DE |
154 |
173 |
30 |
43 |
80.7 |
72.3 |
156 |
223 |
45 |
56 |
77.0 |
67.1 |
224 |
198 |
17 |
65 |
83.9 |
75.0 |
55 |
87 |
17 |
27 |
76.8 |
64.5 |
AR |
101 |
82 |
14 |
203 |
73.8 |
73.5 |
88 |
109 |
19 |
264 |
68.0 |
64.6 |
137 |
77 |
13 |
277 |
84.1 |
78.2 |
29 |
26 |
9 |
122 |
79.6 |
69.4 |
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
idx |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
| GPT-4V |
0 |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
2 |
158 |
181 |
32 |
29 |
84.1 |
81.3 |
163 |
219 |
75 |
23 |
85.1 |
80.8 |
205 |
234 |
34 |
31 |
82.1 |
81.2 |
55 |
95 |
22 |
14 |
77.8 |
75.3 |
5 |
143 |
202 |
31 |
24 |
81.2 |
76.3 |
175 |
216 |
72 |
17 |
88.0 |
81.5 |
179 |
250 |
44 |
31 |
78.6 |
75.6 |
51 |
106 |
18 |
11 |
76.4 |
71.5 |
| LLaVA |
0 |
286 |
98 |
3 |
13 |
54.4 |
52.3 |
331 |
134 |
13 |
2 |
69.7 |
67.7 |
363 |
112 |
28 |
1 |
61.5 |
58.1 |
116 |
65 |
2 |
3 |
56.3 |
55.9 |
1 |
197 |
163 |
7 |
33 |
48.6 |
48.5 |
169 |
236 |
15 |
60 |
69.6 |
66.3 |
247 |
211 |
9 |
37 |
57.6 |
57.9 |
68 |
105 |
5 |
8 |
56.6 |
57.5 |
2 |
170 |
223 |
4 |
3 |
49.1 |
49.5 |
197 |
274 |
7 |
2 |
69.0 |
68.5 |
318 |
178 |
7 |
1 |
58.1 |
57.9 |
41 |
141 |
4 |
0 |
56.0 |
57.0 |
3 |
243 |
145 |
8 |
4 |
52.6 |
53.0 |
243 |
222 |
13 |
2 |
71.8 |
70.2 |
295 |
197 |
11 |
1 |
57.3 |
56.3 |
82 |
93 |
10 |
1 |
56.5 |
56.5 |
4 |
211 |
180 |
6 |
3 |
62.1 |
62.0 |
232 |
244 |
3 |
1 |
77.3 |
77.1 |
210 |
286 |
4 |
4 |
57.3 |
56.3 |
55 |
129 |
2 |
0 |
57.6 |
57.0 |
5 |
241 |
154 |
3 |
2 |
63.3 |
63.0 |
313 |
163 |
4 |
0 |
80.5 |
80.4 |
260 |
236 |
8 |
0 |
62.1 |
62.3 |
109 |
75 |
2 |
0 |
52.7 |
52.7 |
4.3 RQ3: Robustness to Prompts
Multilingual prompts
Since our datasets primarily originate in English-language websites, multilingual prompts can assess the adaptability of intrinsic knowledge across different languages and mitigate data leakage. We translate our prompt template and claims into Chinese, German, and Arabic using the Google Translate API. Table 3 presents the performance across different languages. Notable variations can be observed in adherence to instructions, particularly in Arabic, which exhibits a higher frequency of Deny responses across datasets. Nevertheless, the overall accuracy levels remain comparable, suggesting that the responses’ lack of explanations and confidence scores in Arabic did not significantly impact model performance. Moreover, the True/False Accuracy generally surpasses the Overall Accuracy, indicating GPT-4V’s good knowledge awareness. We do not conduct experiments on open-source models as they are predominantly trained and tested in English.
Prompts with similar semantics
We employ ChatGPT to generate semantically similar yet distinct prompts, as depicted in Figure 3. The part pertaining to explanations and confidence scores is identical to the original prompt. We randomly select two prompts for testing GPT-4V and all generated prompts for the LLaVA(13b) model. The outcomes of the experiments are detailed in Table 4, where Index 0 denotes the original prompt, and the other indices represent the newly crafted prompts. The data reveals that GPT-4V maintains commendable consistency, while LLaVA demonstrates variable biases across different prompts. For the same dataset, such as Fauxtography, models tend to predict the majority of claims as True when using the original prompt (index 0). In contrast, the prompt of index 2 leads the models to predict most claims as False. Moreover, on Fauxtography and COSMOS, the maximum discrepancy in overall accuracy reached and , respectively, compared to just and for GPT-4V. It indicates that the LLaVA(13b) model is highly sensitive to prompts and fails to maintain consistency.
4.4 RQ4: Improving Capabilities via In-Context Learning
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
Demo |
Num |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
0 |
0 |
286 |
98 |
16 |
0 |
54.4 |
52.8 |
331 |
134 |
14 |
1 |
69.7 |
67.7 |
363 |
112 |
29 |
0 |
61.5 |
58.1 |
116 |
65 |
5 |
0 |
56.4 |
55.9 |
| 1 |
1 |
244 |
135 |
9 |
12 |
60.9 |
61.8 |
295 |
171 |
7 |
7 |
75.3 |
74.2 |
255 |
229 |
12 |
8 |
62.6 |
62.7 |
95 |
88 |
3 |
0 |
56.3 |
55.4 |
2 |
182 |
161 |
6 |
51 |
59.5 |
60.0 |
234 |
220 |
8 |
18 |
75.8 |
73.5 |
169 |
282 |
10 |
43 |
57.4 |
55.5 |
42 |
109 |
6 |
29 |
67.5 |
63.4 |
| 2 |
1 |
264 |
121 |
4 |
11 |
61.6 |
62.8 |
337 |
131 |
7 |
5 |
73.9 |
73.5 |
314 |
180 |
4 |
6 |
64.8 |
65.1 |
97 |
79 |
10 |
0 |
57.4 |
55.9 |
2 |
176 |
204 |
10 |
10 |
60.0 |
60.5 |
233 |
230 |
11 |
6 |
80.3 |
79.6 |
199 |
291 |
8 |
6 |
61.2 |
60.7 |
44 |
131 |
11 |
0 |
60.0 |
60.8 |
| 3 |
1 |
211 |
185 |
1 |
3 |
62.6 |
62.8 |
279 |
195 |
6 |
0 |
72.2 |
71.7 |
236 |
266 |
2 |
0 |
61.8 |
61.5 |
75 |
111 |
0 |
0 |
60.2 |
60.2 |
2 |
191 |
198 |
9 |
2 |
64.1 |
63.8 |
263 |
201 |
16 |
0 |
75.4 |
75.4 |
206 |
288 |
9 |
1 |
61.1 |
61.7 |
54 |
127 |
4 |
1 |
59.7 |
59.7 |
| 4 |
1 |
191 |
193 |
9 |
7 |
62.2 |
62.0 |
260 |
203 |
16 |
1 |
78.0 |
77.9 |
210 |
283 |
11 |
0 |
60.6 |
60.9 |
50 |
126 |
8 |
2 |
60.8 |
60.8 |
2 |
196 |
188 |
7 |
9 |
62.8 |
62.8 |
259 |
207 |
12 |
2 |
77.5 |
77.5 |
207 |
284 |
11 |
2 |
61.1 |
61.3 |
52 |
122 |
10 |
2 |
61.5 |
61.8 |
As demonstrated in the 1st and 4th examples in Figure 2, the LLaVA(13b) model struggles to distinguish claims associated with manipulated images. The image content is consistent with the textual claim but contradicts real-world facts. To enhance the model’s comprehension of the task, checking facts rather than the consistency between the image and text, we experiment with employing in-context learning to improve its performance. In detail, we obtain GPT-4V’s responses for claims outside our test datasets and employ these as instructive examples for LLaVA(13b). Due to LLaVA’s limitation with multiple image inputs, we combine different images vertically with a 20-pixel gap between them. The linkage between images and claims is narratively indicated, as illustrated in the following prompt:
@@ 4.4 RQ4: Improving Capabilities via In-Context Learning
We construct four sets of demonstrations, each comprising one true and one false claim. We report the performances when using only the false claim and employing both claims as examples. We observe a universal improvement in model accuracy when examples are presented in Table 5. However, the use of the second example does not contribute to accuracy. In the explanations provided by LLaVA, we can observe two effects: one is the average length of the explanation. For instance, on MOCHEG, with the presentation of 0, 1, and 2 examples from the first set, the average sentence length of the model’s explanations are 3.09, 4.02, and 3.75, respectively. Another change is the frequency of the term "manipulated," which increases from 4 to 170 and 193 on MOCHEG. This suggests that the model can learn to analyze from the perspective of image manipulation. The last two rows of Figure 2 display responses from the LLaVA(13b) model. In the 4th example, it is evident that through learning, the model recognizes the altered part in the image and explains the motive behind the modifications: "but the gun and menacing pose are added elements," and "image is likely created to spread misinformation and discredit Joe Biden or to provoke a reaction from viewers." However, some samples are not correctly identified, such as the 1st example, and we observe that LLaVA produces outdated information, as highlighted in green.
4.5 RQ5: Calibration Evaluation

### v2

@@ 4 Experimental Setups
We use OpenAI’s API to collect responses from GPT-4V (gpt-4-1106-preview). In addition, we experiment with open-source MLLMs, including LLaVA-1.5 (7b and 13b, Liu et al. 2023b, a), and MiniGPT (v2, Chen et al. 2023). All parameters are set to the default values, with max_tokens=300 for GPT-4V. These models are notably representative and competitive in a multimodal setup Fu et al. (2023); Liu et al. (2023c). We further explore two approaches that do not require fine-tuning to enhancing the model’s performance:
Prompt Ensembles (PE)
involve using a variety of prompts for the same task and aggregating model’s responses to produce more accurate and less biased predictions. We employ ChatGPT to generate 5 more semantically similar yet distinct prompts, as depicted in Figure 2. The part about explanations and confidence scores is identical to the original prompt. With these prompts we collect six responses for each claim. Then, we conduct a vote to determine the final response type. If the two highest-scoring response types receive the same number of votes, we consider the model uncertain.
In-Context Learning (ICL)
enables a model to learn tasks by directly observing examples within the prompt, bypassing the need for prior explicit training. Debunking articles, which are lengthy and often contain irrelevant descriptions, are not used here. Instead, we utilize responses from GPT-4V to claims not included in our test datasets as instructive examples. To overcome LLaVA’s limitation in handling multiple image inputs, we vertically stack different images with a 20-pixel separation. The relationship between each image and its corresponding claim is explicitly described, such as "For the first picture, claim: ", as shown in the prompt below. We perform 1 and 2-shot learning, showing the model one and two examples, respectively, referred to as ICL-1 and ICL-2. We use four pairs of examples as demonstrations. For ICL-1, all examples are False, while for ICL-2, the first claim is always False, and the second is True. They are detailed illustrated in Figure 10 in Appendix.
0.5cm0.5cm
@@
5 Experimental Results
We comprehensively analyze the data we collected in Section 5.1. In Section 5.2, we provide a detailed analysis of the model explanations, focusing on their step-by-step reasoning, and discuss our taxonomy of failure reasons. Section 5.3 is dedicated to ablation studies.
5.1 Comprehensive Results
The comprehensive results are displayed in Table 2. The first four columns for each dataset represent the number of different types of responses, while the following are True&False and Overall Accuracy, respectively. We have highlighted the highest metrics and the highest ones among the open-source models. Hereafter, we omit the version numbers of the open-source models and refer to them simply as LLaVA(7b), LLaVA(13b), and MiniGPT. Both PE and ICL experiments are based on LLaVA(13b). LLaVA+PE involves voting based on 6 responses; if the counts of True and False are equal, then it is an uncertain case. LLaVA+ICL-1 and LLaVA+ICL-2 represent the average results across four sets of demonstrations, we have rounded the number of different responses. Detailed outcomes for PE and ICL are presented in Table 5 and 6 in Appendix.
Overall, GPT-4V demonstrates surprising accuracy, with both accuracy metrics exceeding 80% on the three public datasets. The accuracy decrease is not particularly significant on Post-4V, with the True&False Accuracy equals . In addition, these values are mostly higher than Overall Accuracy, suggesting that GPT-4V exhibits higher precision when it responds with confidence. There are more False cases than True ones. The high number of Others is primarily due to GPT-4V safety alignment. This results in frequent replies such as "I’m sorry, but I cannot assist with this request." or "This content may violate our content policy." Additionally, GPT-4V expresses uncertainty the most.
In contrast, MiniGPT cannot provide explanations and confidence, while LLaVA(7b) shows a strong bias, almost always responding True with this prompt. LLaVA(13b) shows improved accuracy across all datasets, better adherence to instructions, and more frequent expressions of uncertainty. We observe that the two approaches, PE and ICL, can enhance the model’s performance to varying degrees. As to PE, there is an obvious improvement in the COSMOS dataset, with increases of 6.6 and 4.0 in two accuracy metrics, respectively. However, the enhancement on other datasets is very limited. ICL shows a greater impact. On the COSMOS dataset, LLaVA+ICL-1 brings increases of 7.6 and 9.0 in both metrics. On MOCHEG, the improvement is minimal, with gains of 0.9 and 4.4, respectively.
Calibration